OpenStudy (anonymous):

Solve for R: 3=(1+(r/12) ) ^(12*21)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

252 ln (1+(r/12) = ln 3 ln (1+(r/12) = ln 3 / 252 = 4.35957257 X 10^-3 1+r/12 = 1.004369089 r/12 = 0.004369089 r = 0.05243 to 5 placeso f deci,als

OpenStudy (anonymous):

Thanks- I got the same thing now so you're hopefully right!

OpenStudy (anonymous):

ok